Joan is three times as old as Anne. Fifteen years from now Joan will be twice as old as Anne then what is their ages now?
Answer by stanbon(75887)

Joan is three times as old as Anne. Fifteen years from now Joan will be twice as old as Anne then what is their ages now?
-------------------
Equations:
J = 3*a
j+15 = 2(a+15)
---------------
Substitute for "j" and solve for "a"::
3a+15 = 2a+30
a = 15 (Anne's age now)
------------
j = 3a = 3*15 = 45 (Joan's age now)
---------
Cheers
Stan H.
----------